OpenStudy (anonymous):

anti-derivative (1)/x^2dx

OpenStudy (amistre64):

\[\int x^{-2}\ dx\]

OpenStudy (amistre64):

to derive a power we multiply and subtract 1, to undo that we add 1 and divide

OpenStudy (anonymous):

So would my answer be (-1/3x^3)?

OpenStudy (amistre64):

it would; +C

OpenStudy (amistre64):

opps, read it wrong: -1/3 x^-3

OpenStudy (amistre64):

+C ....

OpenStudy (amistre64):

\[\int x^{-2}\ dx=-\frac{1}{1}x^{-1}=-\frac{1}{x}+C\] thats better

OpenStudy (anonymous):

okay thanks. I need to know how to type another problem but I do not know how to type it in integral notation. COuld you explain how to do that please?

OpenStudy (amistre64):

I gotta get to class at the moment; but good luck :)

OpenStudy (anonymous):

ok well later Thanks. :)

OpenStudy (anonymous):

use the power rule backwards

OpenStudy (anonymous):

\[\frac{d}{dx}x^n=nx^{n-1}\] \[\int x^n dx=\frac{x^{n+1}}{n+1}\]